is a triangle with sides of 6,8,10,a right triangle
Answer by stanbon(75887)

is a triangle with sides of 6,8,10,a right triangle
------------
Question:
Does 10^2 = 6^2 + 8^2 ??
---
100 = 36+64
100 = 100
-----------
Yes, it is a right triangle.
Cheers,
Stan H.